# FALCON: Resolving Visual Redundancy and Fragmentation in High-resolution Multimodal Large Language Models via Visual Registers

### v1

@@ 4.2 Ablation Study
We perform extensive ablation studies to validate the effectiveness of the proposed ReCompact and ReAtten. All experiments are conducted with 64 visual registers with up to 9 sub-images. In stage 3, a subset of 800k samples is used.
@@
Effectiveness of ReCompact mechanism. To verify the advantages of the proposed ReCompact mechanism in reducing visual redundancy and compressing visual representations during the visual encoding phase, we compare the ReCompact with other commonly used compression methods. This includes pooling layers [60, 48] and the Abstractor [65]. The latter serves as a representative module of the query token combined with cross-attention.
As shown in Fig. 4, ReCompact demonstrates superior performance compared to both pooling and Abstractor. These results indicate that ReCompact effectively eliminates visual redundancy and produces a more expressive compact visual representation. In contrast, we argue that the pooling layer merely reduces the number of tokens by merging adjacent ones, failing to eliminate redundancy that interferes with the LLM’s understanding of visual content. Additionally, although Abstractor uses significantly more parameters, its performance is generally lower than that of pooling and ReCompact. This result suggests that even millions of samples may be insufficient for fully training Abstractor, which diminishes their practical efficacy.
Table 3: Evaluation on benchmarks across diverse domains. The POPE, SQA, and MMBench benchmarks are used to assess hallucination, scientific reasoning, and overall abilities, respectively. MMBench results are taken from the OpenCompass Leaderboard, with * indicating results sourced from respective papers.
Model SQA POPE MMB
@@
Qwen-VL [2] 67.1 - 41.6
InstructBLIP [8] 63.1 78.9 38.4
LlaVA-v1.5-7B [31] 66.8 85.9 62.8
mPLUG-Owl2 [65] 68.7 85.8 63.5
@@
Monkey [28] 69.4 - 57.4
LlaVA-FlexAttn [24] - 85.9 65.7*
S2-Wrapper-7B [48] - - 66.2*
LLaVA-Next-Vicuna7B [30] 70.1 86.5 66.5
SliME-7B [69] 76.8 85.4 63.1
FALCON-3B 84.4 87.0 73.3
FALCON-7B 85.5 87.3 73.6
@@
Figure 6: Ablation on different visual register counts. Increasing the number of visual registers generally enhances accuracy across both benchmarks. However, the improvement rate noticeably tapers off when the register count exceeds 64. Based on this observation, to better balance the efficiency and performance, we select register counts of 64 and 36 for FALCON-7B and 3B, respectively.
@@
Ablation on the Number of Visual Registers. To investigate the impact of changing the number of visual registers on model performance, we conduct experiments with visual register counts set to 36, 64, and 144. These counts correspond to visual token compression rates of 16×16$\times$16 ×, 9×9$\times$9 ×, and 4×4$\times$4 ×, respectively. As shown in Fig. 6, increasing the number of visual registers generally enhances the model’s high-resolution comprehension capabilities and helps reduce hallucinations. However, when the register count exceeds 64, the improvement gains start tapering off noticeably. Meanwhile, increasing visual registers also demands more computational resources. Compared to the baseline model with 36 registers, models configured with 64 and 144 registers require 1.15 and 2.41 times more training time, respectively. Base on these observations, to achieve a better balance between efficiency and performance, we select 64 and 36 as the register counts for FALCON-7B and 3B, respectively.
@@ 4.3 Qualitative Analysis
Visualization and Analysis of Attention Maps. To qualitatively analyze the proposed visual register technique, we visualize the register-to-image attention maps within ViT. Specifically, for a particular visual register, we extract the attention scores that this register assigns to image tokens in each sub-image. We then concatenate these attention scores across all sub-images, create a heatmap, and overlay it onto the input image. As shown in Fig. 5(b), each visual register focuses on relatively distinct areas of the image, highlighting that it captures a diverse range of visual details. Moreover, the visual registers tend to avoid background areas in the image, indicating their effectiveness in identifying and discarding visual redundancy. These properties enable visual registers to effectively produce a comprehensive and non-redundant visual representation. Additionally, to qualitatively evaluate the effectiveness of the ReAtten module, we visualize the attention maps of several particular registers within ViT models with and without ReAtten. As shown in Fig. 5(c), in the ViT model without ReAtten, the attention patterns across different sub-images appear highly fragmented. In contrast, the ViT model with ReAtten exhibits continuous attention patterns, indicating effective information exchange between sub-images.
@@ 5 Conclusion
Supplementary Material
Appendix A Implementation Details
A.1 Architecture
We implement FALCON on pretrained LLMs with two different sizes, including Vicuna-v1.5-7B [7] and Llama-3.2-3B-Instruct [1], resulting in FALCON-7B and FALCON-3B, respectively. Both versions employ the pretrained CLIP-ViT-L/14-336px [45] as vision encoder and integrate a two-layer MLP module with GeLU [12] activation to serve as the vision-language projection layer. For FALCON-7B, the number of visual registers is set to 64, while for FALCON-3B, it is set to 36. These settings result in remarkable visual token compression rates of 9x and 16x, respectively. The maximum number of sub-images is set to 16 for both FALCON-7B and FALCON-3B.
A.2 Training Settings
As described in Sec. 3.3, the training process of FALCON involves three stages. In Stage 1, we train only the visual registers and the MLP using the LLaVA-pretrain-558k [29] dataset. The batch size and initial learning rate are set to 256 and 1e-3, respectively, and training is conducted for 1 epoch. In Stage 2, we unfreeze all parameters and train with the ShareGPT4V-PT [6] dataset along with a subset of the DocStruct4M [14] dataset. The combined dataset includes 1.2M image captions and 600K full-text OCR samples. During this stage, we set the batch size to 256 and the initial learning rate to 2e-5, training the model for a single epoch. In Stage 3, we freeze both the vision encoder and the LLM, and fine-tune the LLM with LoRA [15]. The training dataset contains 1M samples collected from publicly available datasets, with details provided in Tab. 5. This stage employs a batch size of 128, with an initial learning rate set to 2e-4 for LoRA parameters and 2e-5 for other trainable parameters. The training is conducted for one epoch. Across all stages, the AdamW optimizer is employed with a weight decay of 0. Additionally, a learning rate warm-up is applied over the initial steps, representing 3% of the total training steps. We utilize four 80GB A800 GPUs for training. The total training time for FALCON-7B and FALCON-3B was approximately 66 hours and 30 hours, respectively.
Table 5: Instruction-tuning datasets used in stage 3.
Task Datasets Size
Dialogue LLaVA [31], ALLaVA [4], ShareGPT 301K
General VQA VQAv2 [11], OKVQA [39], A-OKVQA [46], GQA [18] 230K
Document ChartQA [40], DocVQA [41], TextVQA [50], TextCaps [49], InfoVQA [42], VisualMRC [52], DeepForm [51], OCRVQA [43], WikiTableQuestions [44] 206K
Grounding RefCOCO [19], Visual Genome [22] 134K
Science AI2D [20], IconQA [36], ScienceQA [37] 46K
Caption ShareGPT4V-Cap100K [6] 50K
OCR SynthDog-en [21] 40K
A.3 Instruction-tuning Datasets
As shown in Tab. 5, we present details of the instruction-tuning dataset used by FALCON in stage 3. This dataset encompasses a diverse range of task types, including dialogue, general VQA, document VQA, grounding, science VQA, caption, and OCR, totaling approximately 1M samples. All data are collected from publicly available dataset.
Figure 9: More visualization results of Register-to-Image Attention. The results in (b) demonstrate that ReCompact effectively eliminates redundancy and captures rich visual details across different image types, including natural scenes and documents. The results in (c) further indicate that the ReAtten module ensures continuity in encoding for various image types.
Appendix B Experimental Details
We perform extensive ablation studies to validate the effectiveness of the proposed ReCompact mechanism and ReAtten module. By default, all ablation experiments use 64 visual registers and up to 9 sub-images. The stage 3 training uses a subset of about 800k samples. The remaining dataset settings and hyperparameters are the same as those described in Sec. A.2.
B.1 Details of the Ablation on ReCompact
Here, we present the implementation details of the ablation study on the ReCompact mechanism. In this study, we remove the visual register and instead use the outputs corresponding to image tokens as the visual representation. We then compress this visual representation through either the pooling module or the Abstractor [65] module before feeding it into the LMM. To ensure fairness, we set the visual token compression rate to 9×9$\times$9 × for both experiments, aligning with the compression rate used in ReCompact. Specifically, in the pooling module, we applied a $3\times 3$ average pooling layer, followed by a two-layer MLP. For the Abstractor, we set the number of query tokens to 64, while keeping other configurations consistent with mPLUG-Owl2 [65]. All methods are trained using the three-stage pipeline, with the same dataset and training hyperparameters.
B.2 Details of the Ablation on ReAtten
Here we provide implementation details for the ablation study on the ReAtten module. For the Shifted Window Attention (W-Atten) [34] module used in TextMonkey [32], the original design places W-Atten in layers 2, 6, 24, 46 of a 48-layer ViT model. To retain a similar relative depth in the 24-layer CLIP-ViT-L/14 model, we inserted W-Atten in layers 2, 3, 12, 23, respectively. All other configurations are consistent with those in TextMonkey. For the Complementary Image Pyramid (CIP) Strategy proposed by MiniMonkey [16], to maintain a total maximum sub-image count consistent with other methods (9 sub-images), we set the maximum number of sub-images to 6 for CIP’s detailed group and 3 for the adaptive group, respectively. All methods are trained using the three-stage pipeline, with the same dataset and training hyperparameters. We initialize from the same checkpoint obtained after the stage 2 training. Then, we apply the different methods mentioned above for stage 3 training on the same instruction tuning dataset.
Appendix C More Visualization of Attention Maps
As shown in Fig. 9, we perform further visualization analysis on the models that incorporate the ReCompact and ReAtten techniques. The results in (b) show that the ReCompact mechanism effectively reduces redundancy while capturing rich visual details across different image types, including natural scenes and documents. The results in (c) further highlight that the ReAtten module maintains continuity in encoding across various image types.
Appendix D Broader Impact and Potential Risk
FALCON employs readily available LLMs, which means it inherently shares some of their limitations. This includes the potential for hallucinating ungrounded text or producing biased results. We mitigate such limitations by performing vision-language instruction tuning on a diverse set of high-quality datasets. Nevertheless, we recommend conducting a thorough investigation into its safety and fairness for the intended use before applying it in practice.

### v2

@@ 4.2 Ablation Study
We perform extensive ablation studies to validate the effectiveness of the proposed ReCompact and ReAtten.
@@
Ablation on Training Data. To eliminate the influence of training data and demonstrate the advantages of the proposed method, we utilize exactly the same dataset as LLaVA-v1.5 [30] and employ a two-stage pipeline for training. As shown in Tab. 3, the proposed method exhibits significant performance advantages across multiple benchmarks when trained on the same dataset.
Effectiveness of ReCompact mechanism. To verify the advantages of the proposed ReCompact mechanism in reducing visual redundancy and compressing visual representations during the visual encoding, we compare the ReCompact with other commonly used compression methods. This includes pooling [54, 47], Pixel Shuffle [7] and Abstractor [59], where the Abstractor serves as a representative module that combines query tokens with cross-attention.
As shown in Fig. 4, ReCompact demonstrates superior performance compared to other compression methods. These results indicate that ReCompact effectively eliminates visual redundancy and produces a more expressive compact visual representation. In contrast, we argue that the pooling layer and Pixel Shuffle merely reduces the number of tokens by merging adjacent ones, failing to eliminate redundancy that interferes with the LLM’s understanding of visual content. Additionally, although Abstractor uses more parameters, its performance is generally lower than that of pooling and ReCompact. This result suggests that even millions of samples may be insufficient for fully training Abstractor, which diminishes their practical efficacy.
Table 4: Evaluation on benchmarks across diverse domains. The POPE, SQA, MMBench and SEED-Bench are used to assess hallucination, scientific reasoning, and overall abilities, respectively.
Model SQA POPE MMB SEED-Img
@@
Qwen-VL [1] 67.1 - 41.6 62.3
InstructBLIP [9] 63.1 78.9 38.4 58.8
LlaVA-v1.5-7B [30] 66.8 85.9 62.8 65.8
mPLUG-Owl2 [59] 68.7 85.8 63.5 64.5
@@
Monkey [28] 69.4 83.7 57.4 64.3
LlaVA-FlexAttn [23] - 85.9 65.7 -
S2-Wrapper-7B [47] - - 66.2 67.9
LLaVA-Next-Vicuna7B [31] 70.1 86.5 66.5 69.6
SliME-7B [66] 76.8 85.4 63.1 67.0
FALCON 88.8 88.0 76.1 73.3
@@
Figure 6: Ablation on different visual register counts. Increasing the number of visual registers generally enhances accuracy across both benchmarks. However, the improvement rate noticeably tapers off when the register count exceeds 64. Based on this observation, to better balance the efficiency and performance, we select register counts of 64 for FALCON.
@@
Ablation on the Number of Visual Registers. To investigate the impact of changing the number of visual registers on model performance, we conduct experiments with visual register counts set to 36, 64, and 144. These counts correspond to visual token compression rates of 16×16$\times$16 ×, 9×9$\times$9 ×, and 4×4$\times$4 ×, respectively. As shown in Fig. 6, increasing the number of visual registers generally enhances the model’s high-resolution comprehension capabilities and helps reduce hallucinations. However, when the register count exceeds 64, the improvement gains start tapering off noticeably. Meanwhile, increasing visual registers also demands more computational resources. Compared to the baseline model with 36 registers, models configured with 64 and 144 registers require 1.15 and 2.41 times more training time, respectively. Base on these observations, to achieve a better balance between efficiency and performance, we select 64 as the register counts for FALCON.
@@ 4.3 Qualitative Analysis
Visualization and Analysis of Attention Maps. To qualitatively analyze the proposed visual register technique, we visualize the register-to-image attention maps within ViT. Specifically, for a particular visual register, we extract the attention scores that this register assigns to image tokens in each sub-image. We then concatenate these attention scores across all sub-images, create a heatmap, and overlay it onto the input image. As shown in Fig. 5(b), each visual register focuses on relatively distinct areas of the image, highlighting that it captures a diverse range of visual details. Moreover, the visual registers tend to avoid background areas, indicating their effectiveness in identifying and discarding visual redundancy. These properties enable visual registers to effectively produce a comprehensive and non-redundant visual representation. Additionally, to qualitatively evaluate the effectiveness of the ReAtten module, we visualize the attention maps of several particular registers within ViT models with and without ReAtten. As shown in Fig. 5(c), in the ViT model without ReAtten, the attention patterns across different sub-images appear highly fragmented. In contrast, the ViT model with ReAtten exhibits continuous attention patterns, indicating effective information exchange between sub-images.